PS/SPS User Meeting
PS-SPS Users Meeting for Week 37 held on September 14th, 2023
The agenda for this meeting is based upon user schedule 2.0.0, please cf. here. Please note the current injector schedule which contains information about MDs, Technical stops, etc. and can be found here.
Next user meeting:
For week 38 on Thursday, September 21st, 2023 10:30 CEST
Meeting room: 874/1-011 (CERN)
News from the PS & SPS Physics Coordinator (E. B. Holzer, M. Jäkel)
PS/SPS Physics Coordination:
September 13th to 14th: Martin Jaekel (as back-up: E. Barbara Holzer)
News from the Facilities Operations Meeting (FOM) (E.B. Holzer)
Check the Page 1 on the machine vistars for most up-to-date information of the beam availability:
https://op-webtools.web.cern.ch/vistar/vistars.php?usr=SPS1
https://op-webtools.web.cern.ch/vistar/vistars.php?usr=CPS
Now, during the LHC ion commissioning period more last-minute changes can happend than usually during the run.
No SPS user beams on:
Tuesday September 12th in the afternoon 14:00 to 20:00
Wednesday 8:00 to 20:00
All proton beam (PS, PSB and AD users) will be stopped Wednesday September 13 in the afternoon 13:30 to 17:30
Injector Schedule
An updated version of the injector schedule for 2023 has been presented at the IEFC Friday September 1st:
- Dedicated Ion commissioning periods (i.e. no beam for north area user) from 8:00 to 20:00: September 5, 6, 12, 13, 20 and 25
- Some of these commissioning periods have already be shortened. Other changes to be expected --> proton beam to NA users might be available
- NA users of week 38 have one additional day of beam time
- Proton beam to NA all day on Wednesday 27th (no MD)
- The proton beam will stop on Thursday 28th Septmber at 8:00
Preview of MDs:
Week 37 :
Proposal to allow short parallel MDs (if really parallel) instead of LHC MD prep and LHC MDs and where there are “empty” days
Next weeks:
- Next dedicated SPS MD during the ion run; moved to week 43 to test wirescanners with high intensity and high duty cycle.
- Long parallel MDs during the ion Run so far planned for ion MDs → Could move to Monday if better for Physics.
- Dedicated Linac4 and PSB high intensity in week 44
Intensity split and wobbling
- Intensities:
T4 target has been changed back to 40mm on Wednesday 13/09.
T2 = 50
T4 = 37 (40mm target, regular hadron beam) / target values is T10 = 22
T6 = 50
PS Machine Report (A. Lasheen)
- Very good availability over the past week (96%), no major fault in the PS!
- Active work to improve transmission for EAST and SPS/LHC ion cycles.
- Beam stop on Wednesday for L4-PSB-SPS, some activities in the PS in the background.
- Start of CERN Shielding Benchmark Facility (CSBF) run in EAST.
SPS Machine Report (Verena)
- Continuing LHC ion commissioning.
- Series of long lasting faults is continuing: several issus with injection kicker system this week. Injection kicker 3 dump resistor replaced and injection kicker
- Announcement: LHC high intensity MD Friday
- Announcement: Depending on LHC MD next week, parallel FT ion commissioning
Safety / Radiation Protection (Alex Schouten)
- Helmets in experimental areas
- Doors to experimental halls should be closed
- If a lifting table is used, please notify us so we can arrange for a visit by HSE
- If the roof is leaking due to the rain, please report it to make sure it gets resolved
n_ToF (Nikolas Patronis)
- Genral: Physics programme according to schedule. No major issues. Long beam stop tomorrow to change setup in EAR1 (~8h stop starting at 09h00)
- EAR1: 12C(n,cp) reaction study succesfully realized for the low neutron enrgies <20 MeV using SADR. Till Friday morning the same reactin will be studied using DDX setup (up to 200 MeV). On Friday and up to the end of the 2023 physics run transmission setup tests will take place.
- EAR2: The 26Al(n,p/a) reaction study is running. This measurement is aiming to improve our understanding on the chemical element factories in our galaxy following the “snapshot” provided by the INTEGRAL mission of ESA.
- NEAR: SiC detector radiation hardness study finalized this week
- PoT: 1.79E19
East Area Beam Status (N. Charitonidis)
T09: Beam to be checked with VLAST on Thursday. Muon file available for Wednesday night and user magnet installed and operating in DC mode with the power converter of DHZ045 (last corrector magnet of T9 which is therefore disconnected).
T10: Occasional issues with magnet BHZ027, otherwise good operation.
T11: Shutter open in preparation for run next week.
East Area Users Tour de Table
T8 Main: IRRAD/CHARM & CHIMERA MD (F. Ravotti / S. Fiore & E. Johnson/ R. Froeschl)
Very good week, no issue with operation. Weekly intensity level close to the target value (~2e16 p/w, see plot below) considering the time spared for the ion MD (see Eliott report) on Sep. 6th and the need to reduce the flux from Tuesday 12th for the need of one IRRAD experiment. Beam alignment and center on both X- and Y-axis (e.g. number of spills centered within +/- 2mm): ~98%, excellent on both axes.
During the access on Wednesday, in IRRAD we exchanged the dosimetry and moved all long-term experiments out of beam before the beginning of the RP week. CMS Pixel setup will restart for a 10h run in one week on Tuesday afternoon. In CHARM we removed the user rack, installed a new one with CMS and BE-CEM setups. This will be a parasitic irradiation during RP week ops .
RP week at CHARM/CSBF started yesterday, together with 4 colleagues from KEK (Japan). Excellent beam quality. Program for this year:
- Active neutron measurements with CsI detector
- Exposure of CR-39 dosemeters to neutron fields
- Activation experiments using HL-LHC beam dump materials (Ti, Steel 318)
- Shielding property measurements of heavy concretes
HEARTS/CHIMERA
- Successful VHE heavy ion beam commissioning to CHARM where the operational conditions from 2022 were retrieved. This beam is now ready for the October ion electronic testing activity that will run from the 17th of October to the 1st of November 2023.
- During the MD of the 6th of September, the NOMINAL intensity from LEIR was taken with 3.3E10 charges. RFKO slow extraction was used to modulate the intensity of extracted beam.
- The energies used during this MD were 0.65, 0.75, 1.0 and 2.0 GeV/u with measurement on the diode (with enhanced energy resolution compared to 2022) and on the SRAMs through a remotely moveable table. The range in which the machine, extraction and transfer line can be linearly scaled is between 0.5 and 2.7 GeV/u and can be easily controlled through the B-field of the PS.
T9 Outgoing Main: RE44 HERD (Nicola Mori)
Data acquisition proceeded smoothly with muons for MIP calibration, electron energy scan for TRD calibration and CALO energy resolution studies, and pions for PSD position scan.
On Monday 11th we dismounted part of our setup and successfully transferred it to the North Area; PSD and TRD remained in T9 until the end of the beam slot, and finally dismounted on Wednesday 13th.
On 13th we mounted in H4 behind CMS-HGCAL, had the safety inspection, and had the chance to take data last night with mixed results (some subsystems needs debugging), but overall we are ready to profit of beam time that CMS-HGAL will eventually grant us.
T9 Main: VLAST (LI Xiang)
The Detector has been properly installed into the T09 beam region, and it works normally. We have get some muon beams for MIP calibration since last night.
There is a problem for the rain leakage from the roof to the beam region.
T9 Incoming Main: ALICE PHOS - Week 38 (Mikhail Ippolitov)
ALICE PHOS in preparation to test. We will measure energy and time resolutions with new photodetectors and continue test of front end card. All needed parts delivered. Solution with broken gate in bldg.167 found. Transport for moving and installation systems requested. Safety inspection confirmed on 17-00 20.09.2023.
T10 Outgoing Main: RE7 GAMMA MEV (Nicola Mazziotta)
Nice week. We started our data taking on Wed Sep. 6 evening with -11 GeV/c beam momentum but soon after there was an issue with a magnet. Then we have decided to run the rest of the week at -10 GeV to not stress that magnet. We have dismounted our set-up yesterday early in the morning to leave the room to BL4S. Thanks to all.
T10 Main: BL4S (Berare Göktürk, Martin Schwinzerl, Sarah Zoechling, Markus Joos)
We have installed to the area yesterday successfully. We will be using one DESY table and two XCSA tables, and we are thankful to Aboubakir et al for installing & testing them. Our safety check is done at 9 a.m. today. We will hopefully start taking data today in the afternoon. Currently doing the physical allignment of the detectors. After some calibration measurements, we plan on testing the low pressure Cherenkov detector (XCET.043) first. Special thanks to Maarten for helping us on many many matters! And our students are arriving today (yey)!
T11 Incoming Main: CLOUD - Week 38 (Please Add Your Name Here)
AWAKE (G. Zevi Della Porta)
Week summary:
- Continued exploring plasma density step configurations
Daily summary:
- Monday: difficult beam conditions (NA Damper issue in the morning, then MD-related instability, then SPS access). Proton-laser timing scans with 3% density step at 1.75m
- Tuesday-Wednesday: no beam (lead commissioning in SPS)
- Thursday: almost no beam (SPS recovering from magnet exchange)
- Friday: Good beam. Proton-laser timing and proton intensity scans with 0%, 3% and 6% density steps at 1.75m
- Saturday: septum current monitor interlock (FMCM) issue solved at 12.30. Then good beam. 0% and 3% steps at 1.75 m
- Sunday: Good beam until 3pm, then frequent losses during ramp. Proton-laser timing scans with 0%, 1.5%, 3%, 4.5% steps at 1.75m
- Monday: SPS losses during ramp, then SPS bunch arrival jitter. Density steps at -2%, 0%, +1.5% at 1.75m, starting from a high density plasma (6.2E14/cm3).
Issues:
- UV laser (for electron gun) incompatible with no-pre-pulse configuration. To be addressed before October run
- Vapor source OTC circuit failure on Thursday, despite another intervention by contractors on Tuesday/Wednesday
- SPS: MSE FMCM, solved by Piquet
- SPS: losses during ramp starting Sunday 3pm, solved Monday
- SPS: bunch arrival jitter on Monday, solved same day
Plans for current week:
- Last day of protons (Monday), then 4 weeks of access: recover UV, fix OTC, refurbish laser, install μ-metal to shield Earth’s B field.
North Area Beam Status
Beamlines H2, H4, H6, H8 (N. Charitonidis)
H2: Smooth operation. A few power converter issues solved by the first line. Moving beam test tentantively planned for week 38.
H4: Smooth operation.
H6: No issues.
H8: Smooth operation.
Beamlines P42-K12 and M2 (N. Charitonidis)
M2: 100 GeV muon beam preliminarily tuned with AMBER. Request for one day of high intensity test of the hadron beam with 100 units on T6 on 19th September. Switch back to 50 units in the evening ~ 18:30.
P42/K12: In general, smooth operation. A few magnet trips occured (bend and quads in P42).
North Area Users Tour de Table
Beamlines P42-K12 and M2
K12-P42 Main: NA62 (Sergey Kholodenko)
During the past week, the availability of the beam was far from ideal.
There are three main (and with similar time impact) contributions to the beam availability on T10 (when T4 is ON):
- Verifying the K12/P42 beamlines. We performed it with Nikos on the phone immediately after getting the beam back. (Three stops in the past week)
- Issue with BEND.043.022. (13/09, ~19h40) Nikos called first line and Piquet solved issue.
- Issue with LKr readout crate, that went off at 00h30 on Sun. Access was open for a couple of hours.
In general, when the beam is present we are smoothly taking data (DAQ eff ~ 94.5% during the past week). Spill structure (and FFT), intensity and other parameters are well inline with our requirements. The plan is to take data till the end of the run with this conditions.
Starting from Fri. 15/09 Alina Kleimenova is becoming NA62 Run Coordinatrice till the end of the run.
M2 Outgoing Main: MUONE (Clara Matteuzzi)
MUonE ended as M2 main user on monday morning at 9am, but for a kind opportunity given us by AMBER, we run with the full apparatus (tracker+calorimeter) also during the night of the 12th september. On the 12th at 8am, calorimeter was dismounted and AMBER took over.
The tracker (3 stations) will stay on the beam line and run parasitically with the agreement of AMBER (which we thank very warmly!)
M2 Main: AMBER (Maxim Alexeev)
AMBER is in the commissioning phase for the second part of the data tacking.
At the moment we have cleared the majority of the safety checks and are preparing the sub-systems.
We plan to:
- Continue the commissioning till Monday 18.09
- Conduct on Monday or Tuesday a High intensity test (100 units on the T6) to investigate the beam line performance in view of the observations done during the first running period
- Be ready to start operation of the TPC with Hydrogen in the second part of the next week.
We would like to express many thanks to all the CERN colleagues that contributed to the successful changeover. A particular gratitude goes to the safety team that followed all the process.
H2
H2 Incoming Main: NA61 SHINE - Week 38 (P. Podlaski)
- during this period we will calibrate our hadron calorimeter (PSD)
- no installation necessary (PSD in place, out of the beam, we will move it in)
- all needed vacuum modifications agreed with Nikos and BE-EA
H2 Main: NA65 (Toranosuke Okumura)
Brief presentation
The data taking progress: ~60%
We hope to finish around 18th - 19th, and then perform some tests.
We want 20 GeV and 30 GeV beam on 18th or 19th, for momentum estimation validation with the DsTau detector.
H4
H4 Outgoing Main: ALICE FOCAL (Yury Melikyan)
Performance study of the prototype for the future ALICE Forward Calorimeter (hadron part only).
Beams used:
Hadrons: 60 GeV – 350 GeV
Electrons: 20 GeV – 250 GeV
Muons: 150 GeV
Studies performed:
Energy resolution measured with the CAEN DT5202 readout system at various SiPM bias;
Similar studies repeated with the VMM/SRS readout to compare the results;
Energy resolution at minor rotation angles wrt the beamline (within 2°).
Summary:
Sufficient beam availability ~130 hours;
Key points of the test beam plan accomplished, thanks to the entire SPS team, special thanks to Nikos and Bastien!
Data analysis gears up, here just very preliminary results of the energy scan and bias scan:
H4 Main: CMS HGCAL (Arnaud Steen)
Installation of our setup went smoothly. Safety clearance visit was done at 16:30.
Starting acquiring data with beam this morning.
Plan for the next days:
- scan the position of the detector with 200 GeV electron beam
- take energy scan with electron beam in 1 position
- remove our absorber plates
- scan the position of the detector with pion beam
- long run with muon beam
- repeat these steps with 4 different modules
Plan to change our readout electronic on Monday 18/09 and re-use the same system as in August. If possible, we may need to go back to the beam files we had in August which had a larger rate than what we are currently using.
H6
H6 Parallel: ATLAS ITK PIXEL (Please Add Your Name Here)
H6 Incoming Parallel: ATLAS BCM PRIME - Week 38 (Please Add Your Name Here)
H6 Main: CMS PIXEL (Please Add Your Name Here)
H6 Incoming Parallel: CMS MTD - Week 38 (Please Add Your Name Here)
H8
H8 Main: UA9 (L.S. Esposito)
- Yesterday, we got the UA9 setup up in PPE128.
- Currently, we’re in the process of commissioning the telescope with the beam. Once that’s done, we’ll kick off the data taking.
- Safety inspection was done yesterday. The safety clearance will be issued after a few pictures are sent to the Safety Office.
- The beam looks fine (about ~1M particles/spill). Thanks to Maarten and Anna (EA).
- Coordination of activities is happening between UA9 (week 37) and STI (week 38).
- Our goal is to measure the first crystal assemblies produced at CERN (STI) and crystals for UA9, specifically for the beam merging technique.
H8 Incoming Parallel: STI - Week 38 (L.S. Esposito)
- There will be no changes in the setup in PPE128 from previous week.
- Data taking will proceed with a collaborative schedule involving STI and UA9.
H8 Parallel: LHCB RICH (L. Martinazzoli)
PPE138
Installation completed, safety visit passed, ready to take data.
H8 Outgoing Main: ATLAS TILECAL (Giulio Usai & Tigran Mkrtchyan)
- Successfully completed the data-taking program
- Beams used: Positrons from 10 to 100 GeV, Hadrons from 10 to 180 GeV, Muons at 160 GeV
- Observed high purity for all types of beams
- Ramping up analysis efforts
- Now back into the garage position in PPE158
We will follow up privately
- with feedback on the beams, as soon as we proceed with analysis (it is great to have again electrons!) .
- and with request for the students EUROLABs founds.
Many thanks to all personel of SPS and NA (in particular Maarteen, Anna, Michael,Sylvain,Barbara,Martin,…)
for their great work and exquisite attention that make this weeks alway successful, pleasant and rewarding.
H8 Parallel: CMS MTD BTL (Giulia Sorrentino)
- Setup installed yesterday, safety clearance issued in the evening
- We are proceeding with alignement this morning, then we’ll move to regular testing of our sensors
H8 Parallel: LHCB SCIFI (Ettore Zaffaroni and Federico Ronchetti)
- Setup was already installed and didn’t change. We resumed data taking yesterday evening.
- The roof of the building leaked water on our control room (887/1-Q67) and some electric equipment. We called the firefighers, which also alerted EN-EL. The problem was solved and there were no consequences, as far as I know.
- We were, albeit for a short time, the only users with a rooftop pool.
H8 Parallel: TOTEM PPS (F. Garcia)
The setup for this test beam is ready for starting data taken.
The DUTs are: two Diamond detector packages installed in a Roman pot station.
In addition, the is the ongoing commissioning of the TOTEM triple GEM reference tracker.
The new gas line was tested and everything is fine.
Safety insepction was succesful done yesterday.
The beam parameters are ok.
The plan is to start aligment during the day and carry out the scans of the each of the Diamond detector.
Many thanks for the support from all crews of the NA and services of the H8 beamline.
Parasitic Users
M2 Outgoing: AMBER (Please Add Your Name Here)
H6 Outgoing: CEDAR N (Please Add Your Name Here)
H6 Outgoing: MONOLITH (Please Add Your Name Here)
H6 Outgoing: EP PIXEL (Please Add Your Name Here)
M2: MUONE (Please Add Your Name Here)
H4: RE44 HERD (Nicola Mori)
Please see the HERD report for PS T9.
H6: ATLAS MALTA (Please Add Your Name Here)
H8: STRAW TRACKER RD (Please Add Your Name Here)
H8: CMS RPC (Please Add Your Name Here)
H6 Incoming: ATLAS AFP TOF - Week 38 (Please Add Your Name Here)
AoB
Minutes by the respective speakers, edited by E. B. Holzer, M. Jaekel, and M. Schwinzerl